[The following information applies to the questions displayed below.]

Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labor-hours and its standard cost card per unit is as follows:

Direct material: 6 pounds at $8.00 per pound $ 48.00
Direct labor: 4 hours at $17 per hour 68.00
Variable overhead: 4 hours at $4 per hour 16.00
Total standard variable cost per unit $ 132.00

The company also established the following cost formulas for its selling expenses:

Fixed Cost per Month Variable Cost per Unit Sold
Advertising $ 370,000
Sales salaries and commissions $ 440,000 $ 29.00
Shipping expenses $ 20.00

The planning budget for March was based on producing and selling 19,000 units. However, during March the company actually produced and sold 24,000 units and incurred the following costs:

  1. Purchased 160,000 pounds of raw materials at a cost of $7.20 per pound. All of this material was used in production.
  2. Direct-laborers worked 72,000 hours at a rate of $18.00 per hour.

  3. Total variable manufacturing overhead for the month was $336,960.

  4. Total advertising, sales salaries and commissions, and shipping expenses were $374,000, $540,000, and $285,000, respectively.

rev: 11_16_2018_QC_CS-146879

11. What is the variable overhead rate variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)

12. What amounts of advertising, sales salaries and commissions, and shipping expenses would be included in the company’s flexible budget for March?

13. What is the spending variance related to advertising? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)

14. What is the spending variance related to sales salaries and commissions? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)

15. What is the spending variance related to shipping expenses? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)
Answer #1
11
Variable overhead rate variance 48960 U =336960-(72000*4)
12
Advertising 370000
Sales salaries and commissions 1136000 =440000+(24000*29)
Shipping expenses 480000 =24000*20
13
Spending variance related to advertising 4000 U =374000-370000
14
Spending variance related to sales salaries and commissions 596000 F =1136000-540000
15
Spending variance related to shipping expenses 195000 F =480000-285000